Determination of Percent KHP in a mixture by Acid/Base Titration


Standardization of NaOH:

Ttrial 1:

  • mass of KHP = 0.500 g
  • buret reading (final) = 25.00 mL
  • buret reading (initial) = 0 mL
  • NaOH used = ?
  • molarity of NaOH: ?M


Trial 2:

  • mass of KHP = 0.500 g
  • buret reading (final) = 25.40 mL
  • buret reading (initial) = 0 mL
  • NaOH used = ?
  • molarity of NaOH: ?M


Trial 3:

  • mass of KHP = 0.500 g
  • buret reading (final) = 24.80 mL
  • buret reading (initial) = 0 mL
  • NaOH used = ?
  • molarity of NaOH: ?M


Average molarity from all 3 trials: ?


Determination of % KHP in unknown sample:


Trail 1:

  • mass of sample: 0.600 g
  • buret reading (final) = 16.00 mL
  • buret reading (initial) = 0
  • NaOH used = ?


Trail 2:

  • mass of sample: 0.600 g
  • buret reading (final) = 15.50 mL
  • buret reading (initial) = 0
  • NaOH used = ?


Trail 3:

  • mass of sample: 0.600 g
  • buret reading (final) = 15.80 mL
  • buret reading (initial) = 0
  • NaOH used = ?


calculation of mass of KHP and % KHP = ?



1
Expert's answer
2021-03-26T05:38:22-0400

"Molarity=mass\/rfm"

Trial 1= 25/40=0.625M

Trial 2= 25.4/40= 0.635M

Trial 3 = 24.8/40=0.62M

Average molarity= 0.63M